Operating profit                       $ 25.0

Bond interest                  18.0

Profit before tax                  7.0

Corporation Tax               2.1

Profit after tax                 4.9

Proposed net dividend for year                       4.2

Retained Profit                                         0.7

The bond have a total nominal value of $ 360 million and have a current market value of 350 million. The ordinary share have a total nominal value of $ 5 million and total market value of 120 million. The public beta for SCM plc is 1.5.The risk free rate of return is 3% gross . The market rate of return is 7% .Corporation tax is 30%.

Required: 1) Calculate SCM’s WACC , using market values.

2)given the recent economic downturn, SCM has some liquidity problems .For each of the following area of the Co. working capital , discuss key recommendations on how they can be managed to improve the liquidity position of the company.

-Debtors and credit control

-stocks

-creditors

-short-term investment of surplus cash.
Answer #1

Answer no. 1 Calculation of WACC based on Market Value

Cost of Ordinary Shares - Formula - Risk free rate of return + (Market rate of return - Risk free rate of return)* Beta

= 3 + (7 - 3) * 1.5

= 9 %

Cost of Bonds -Formula - Rate of Interest on bond *(1 - Tax Rate)

Rate of interest = Interest on Bonds *100 / Nominal value of Bonds

= 18*100 / 360

= 5 %

So, the cost of bonds = 5 * (1 - 0.30)

= 3.5

So WACC = (Cost of ordinary shares * market value of ordinary shares / Total Market value of shares and bonds) + (Cost of bonds * market value of bonds / Total Market value of shares and bonds)

= (9 * 120 / 470) + (3.5 * 350 / 470)

= 2.297 + 2.606

= 4.903 % Answer

Note : Total market value of ordinary shares and bonds = 120 + 350 = 470